You want to have $2.65 million when you retire in 36 years. You feel that you can save $550 per month until you retire. What APR do you have to earn in order to achieve your goal?
Answer #1

Ans 10.51%

Future Value of Annuity = P ( (1 + r)n - 1 ) / r
2650000 = 550* ((1 + r%/12)^432 - 1) / (r%/12)
2650000 / 550 = ((1 + r%/12)^432 - 1) / (r%/12)
4818.18181818182 = ((1 + r%/12)^432 - 1) / (r%/12)
r = 10.51%